# TacCompress: A Benchmark for Multi-Point Tactile Data Compression in Dexterous Manipulation

### v1

@@ IV-D Lossy Compression Evaluation
Fig. 6, Fig. 7, Fig. 8, and Fig. 9 illustrate the rate-distortion performance when applying lossy compression, using PSNR [42] and MS-SSIM [43] as the distortion metrics. Given the similarity of the tactile images to screen content, we first apply the state-of-the-art screen content coding (SCC) tools [44] in HM and VTM, comparing their performance against standard intra coding mode, as shown in Fig. 6 and Fig. 7. Notably, HM and VTM maintain reconstruction quality of no less than 38dB at extremely low bit rates (0.01 bpss, approximately 800 compression ratios) even using normal intra encoding mode, demonstrating their effectiveness for this task. Similar to lossless compression, lightweight and geometrically regular objects, such as eggs and water bottles, achieve better rate-distion performance than objects like apples and vitamin bottles, as shown in Fig. 6. Meanwhile, it can be observed that for the same codec, enabling SCC tools largely improves the compression efficiency of tactile images, achieving 30.67% and 56.42% BD-Rate [45] gains for VTM and HM, respectively.
@@
We also evaluate other lossy image codecs like JPEG2000, JPEG-XL, and TCM, as shown in Fig. 8 and Fig. 9, where the prior demonstrate for distinct objects’ compression performance, the latter shows the average performance across all the objects. The three methods are all general-purpose codecs designed for all types of images and deliver moderate compression ratios for our tactile signals. They exhibit inferior encoding performance compared to HM and VTM in both PSNR and MS-SSIM. Neural-based compressors like TCM and L3C often struggle with the generalization capability and perform sub-optimally on such out-of-domain data. However, while TCM’s PSNR performance is inferior to JPEG2000, it excels in term of MS-SSIM metric.
IV-E Applications on Downstream Tasks
@@
To validate the effectiveness of the proposed dataset for downstream tasks, we conduct object classification experiments on the generated tactile images. First, we perform a clustering analysis to verify the separability of the generated data. The uncompressed images are projected to 2D space using t-SNE [46], followed by K-means clustering [47] with set to the number of object categories. As illustrated in Fig. 10, the tactile images of each object form distinct clusters, indicating that object-specific features are preserved in the dataset and it is feasible to perform object classification.
@@
Then, we simply employ some classical machine learning (ML) methods, including linear regression (LR) [48], support vector machine (SVM) [49], random forest (RF) [50], and K-nearest neighboring (K-NN) [51], to classify the corresponding object of each tactile image. To ensure a balanced evaluation, 70% of the images from each object are randomly selected as the training set, while the remaining serve as the test set. The images are normalized and flattened into vectors as inputs to the classification models. As presented in Table. II, when classifying uncompressed raw images, all these methods achieve at least 70% accuracy, among which SVM and LR perform the best, with accuracies of 81.25% and 79.17%, respectively. LR’s satisfactory performance suggests that tactile image data exhibit relatively well-structured feature distributions. The superior performance of SVM can be attributed to its ability to handle high-dimensional feature spaces and effectively separate complex patterns, making it well-suited for tactile image classification.
Further, we also analyze the impact of compression on object classification using VTM-SCC, the most efficient lossy compressor. As shown in Table. II, compared to performing object classification on raw data, VTM reduces the data volume by a factor of 1200 (i.e., 0.0067 bpss), while still maintaining high classification accuracy (over 70% with SVM and LR). When the data volume is compressed by a factor of 500 (i.e., 0.016 bpss), the resulting distortion has minimal impact on the classification task, particularly for SVM (with only 1% accuracy loss compared to classifying the uncompressed raw data). These results demonstrate the robustness of the proposed dataset for downstream tasks and highlight the effectiveness of our compression methods in reducing tactile data size while retaining essential information for accurate classification.
V Conclusion
Data | bpss | SVM | RF | K-NN | LR |
@@ V Conclusion
0.0110 | 77.08% | 71.88% | 64.58% | 73.96% | |
0.0067 | 72.92% | 66.67% | 61.45% | 70.83% |

### v2

@@ 4.4. Lossy Compression Evaluation
fig. 6, fig. 7, fig. 8, and fig. 9 illustrate the rate-distortion performance when applying lossy compression, using PSNR (Korhonen and You, 2012) and MS-SSIM (Bakurov et al., 2022) as the distortion metrics. As shown in fig. 5, tactile images exhibit several characteristics commonly found in screen content, such as sharp contours, limited color variation, and repetitive spatial patterns. These properties make screen content coding (SCC) tools particularly suitable for compressing such data. Therefore, we apply the SCC tools in HM and VTM (Xu and Liu, 2021), and compare their performance against the standard intra coding modes, as shown in fig. 6 and fig. 7. Notably, HM and VTM maintain reconstruction quality of no less than 38 dB at extremely low bit rates (0.01 bpss, approximately 800 compression ratios) even using normal intra encoding mode, demonstrating their effectiveness for this task. Similar to lossless compression, lightweight and geometrically regular objects, such as eggs and water bottles, achieve better rate-distortion performance than objects like apples and vitamin bottles, as shown in fig. 6. Meanwhile, it can be observed that for the same codec, enabling SCC tools largely improves the compression efficiency of tactile images, achieving 30.67% and 56.42% BD-Rate (Bjontegaard, 2001) gains for VTM and HM, respectively.
@@
We also evaluate other lossy image codecs like JPEG2000, JPEG-XL, and TCM, as shown in fig. 8 and fig. 9, where the prior demonstrates distinct objects’ compression performance, and the latter shows the average performance across all the objects. The three methods are all general-purpose codecs designed for all types of images and deliver moderate compression ratios for our tactile signals. They exhibit inferior encoding performance compared to HM and VTM in both PSNR and MS-SSIM. Neural-based compressors like TCM and L3C often struggle with generalization capability and perform sub-optimally on such out-of-domain data. However, while TCM’s PSNR performance is inferior to JPEG2000, it excels in terms of the MS-SSIM metric.
4.5. Applications on Downstream Tasks
@@
To validate the effectiveness of the proposed dataset for downstream tasks, we conduct object classification experiments on the generated tactile images. First, we perform a clustering analysis to verify the separability of the generated data. The uncompressed images are projected to 2D space using t-SNE (Van der Maaten and Hinton, 2008), followed by K-means clustering (Hartigan and Wong, 1979) with set to the number of object categories. As illustrated in fig. 10, the tactile images of each object form distinct clusters, indicating that object-specific features are preserved in the dataset and it is feasible to perform object classification.
@@
Then, we simply employ some classical machine learning (ML) methods, including linear regression (LR) (Seber and Lee, 2012), support vector machine (SVM) (Hearst et al., 1998), random forest (RF) (Rigatti, 2017), and K-nearest neighboring (K-NN) (Peterson, 2009), to classify the corresponding object of each tactile image. To ensure a balanced evaluation, 70% of the images from each object are randomly selected as the training set, while the remaining serve as the test set. The images are normalized and flattened into vectors as inputs to the classification models. As presented in table 2, when classifying uncompressed raw images, all these methods achieve at least 70% accuracy, among which SVM and LR perform the best, with accuracies of 81.25% and 79.17%, respectively. LR’s satisfactory performance suggests that tactile image data exhibit relatively well-structured feature distributions. The superior performance of SVM can be attributed to its ability to handle high-dimensional feature spaces and effectively separate complex patterns, making it well-suited for tactile image classification.
| Data | bpss | SVM (Hearst et al., 1998) | RF (Rigatti, 2017) | K-NN (Peterson, 2009) | LR (Seber and Lee, 2012) |
@@
| 0.0110 | 77.08% | 71.88% | 64.58% | 73.96% | |
| 0.0067 | 72.92% | 66.67% | 61.45% | 70.83% |
Further, we also analyze the impact of compression on object classification using VTM-SCC, the most efficient lossy compressor. As shown in table 2, compared to performing object classification on raw data, VTM reduces the data volume by a factor of 1200 (i.e., 0.0067 bpss), while still maintaining high classification accuracy (over 70% with SVM and LR). When the data volume is compressed by a factor of 500 (i.e., 0.016 bpss), the resulting distortion has minimal impact on the classification task, particularly for SVM (with only 1% accuracy loss compared to classifying the uncompressed raw data). These results demonstrate the robustness of the proposed dataset for downstream tasks and highlight the effectiveness of our compression methods in reducing tactile data size while retaining essential information for accurate classification.
5. Conclusion
6. Acknowledgment
This work was partly supported by the Paxini Tech Research Fund, NSFC62431015, Science and Technology Commission of Shanghai Municipality No.24511106200, the Fundamental Research Funds for the Central Universities, Shanghai Key Laboratory of Digital Media Processing and Transmission under Grant 22DZ2229005, 111 project BP0719010.